OpenStudy (anonymous):

A bakery sells 6 bagels for $2.99 and 4 muffins for $3.29. What is the total cost in dollars for 4 dozens bagels and 16 muffins, not including tax?

OpenStudy (agent47):

You're buying 4*12=48 bagels and 16 muffins. 6 bagels cost $2.99, and you're buying 8 times that (8*6=48), so you are paying 8*2.99 for bagels.

OpenStudy (agent47):

4 muffins cost $3.29, and you're buying 4 times that (4*4=16), so you are paying 4*3.29 for muffins.

OpenStudy (agent47):

So overall you're paying: 8*2.99+4*3.29 dollars for everything.

OpenStudy (anonymous):

Thank you!